1)Find a unit vector in the direction of the given vectors
a) (3,-4) b)(24,7)
2)Now find unit vector , different from the corresponding one obtained in previous exercise, in the direction of the given vectors
a)(3,-4) b) (24,7)



Answer by stanbon(48564)

1)Find a unit vector in the direction of the given vectors
a) (3,-4)
Find tan^-1(-4/3) = -53.13 degrees
The x-coordinate of the unit vector is cos(-53.13) = 0.6
The y-coordinate of the unit vector is sin(-52.13) = -0.8
b) (24,7)
Find tan^-1(7/24) = 16.26 degrees
The x-coordinate of the unit vector is cos(16.26) = 0.96
The y-coordinate of the unit vector is sin(16.26) = 0.28
===========================



2)Now find unit vector , different from the corresponding one obtained in previous exercise, in the direction of the given vectors
a)(3,-4)
Multiply the coordinates found above by some constant like "2".
(2*0.6,2*(-0.8)) = (1.2,-1.6)
----------------
b) (24,7)
Do the same for the x and y coordinates found above .
Cheers,
Stan H.